OpenStudy (anonymous):

what is the simplest form for 3^√32

OpenStudy (anonymous):

@aajugdar @SandeepReddy

OpenStudy (anonymous):

what is this?

OpenStudy (anonymous):

Can you write it using Equation tab

OpenStudy (anonymous):

the power is root 32?

OpenStudy (anonymous):

okay @SandeepReddy you take it from here m off

OpenStudy (anonymous):

Equation tab is near bottom right of commenting box

OpenStudy (anonymous):

good night (:

OpenStudy (anonymous):

Okay @aajugdar

OpenStudy (anonymous):

@aajugdar you must be indian?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Im too AP

OpenStudy (anonymous):

\[\sqrt[3]{32}\]

OpenStudy (anonymous):

@SandeepReddy

OpenStudy (anonymous):

32 can be written as 2*16 =2*2*8 = 2*2*2*4 =(2^3)*4 right

OpenStudy (anonymous):

uhm ?

OpenStudy (anonymous):

\[\sqrt[a]{x^b} =x ^{b/a}\]

OpenStudy (anonymous):

Do u knw this?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

That is the relation we have in real numbers Remember it!

OpenStudy (anonymous):

ohkay lol

OpenStudy (anonymous):

now 32 can be written as \[32 = 2^5\]

OpenStudy (anonymous):

\[\sqrt[3]{2^5} = 2^{\frac{ 5 }{ 3 }}\]

OpenStudy (anonymous):

ohkay

OpenStudy (anonymous):

so thaats the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks. I have one more, if you dont mind...

OpenStudy (anonymous):

Can be xpressed in so many other ways though!

OpenStudy (anonymous):

Yeah ask!

OpenStudy (anonymous):

\[2\sqrt[3]{4}\]